On December 31, 2017, Dow Steel Corporation had 790,000 shares of common stock and 49,000 shares of 7%, noncumulative, nonconvertible preferred stock issued and outstanding. Dow issued a 4% common stock dividend on May 15 and paid cash dividends of $590,000 and $88,000 to common and preferred shareholders, respectively, on December 15, 2018.

On February 28, 2018, Dow sold 72,000 common shares. Also, as a part of a 2017 agreement for the acquisition of Merrill Cable Company, another 19,000 shares (already adjusted for the stock dividend) are to be issued to former Merrill shareholders on December 31, 2019, if Merrill's 2019 net income is at least $690,000. In 2018, Merrill's net income was $820,000.

In keeping with its long-term share repurchase plan, 6,000 shares were retired on July 1. Dow's net income for the year ended December 31, 2018, was $3,050,000. The income tax rate is 40%.

As part of an incentive compensation plan, Dow granted incentive stock options to division managers at December 31 of the current and each of the previous two years. Each option permits its holder to buy one share of common stock at an exercise price equal to market value at the date of grant and can be exercised one year from that date. Information concerning the number of options granted and common share prices follows:

Date Granted Options Granted Share Price

(adjusted for the stock dividend)

December 31, 2016 26,000 $ 44

December 31, 2017 21,000 $ 53

December 31, 2018 24,500 $ 52

The market price of the common stock averaged $52 per share during 2018.

On July 12, 2016, Dow issued $500,000 of convertible 10% bonds at face value. Each $1,000 bond is convertible into 40 common shares (adjusted for the stock dividend).

Required:

Compute Dow's basic and diluted earnings per share for the year ended December 31, 2018.
Answer #1

Answer:

Numerator / Denominator = Earnings per share
Dow's basic 2,962 / 881 =                3.36
Dow's diluted 2,992 / 924 =                3.24

Calculation

Options which was issued in 2018 does not affect the EPS calculation of 2018. It is because the exercise price is 52 which is equal to the market price. So the options are not dilutive and anti dilutive.

Options which was issued in 2017 is not taken for calculating EPS. It is because the share price is 53 which is higher than average market price of 52. So, these options are antidilutive.

The options issued in 2016 are exercised for 26,000 shares for the calculation of 2021 EPS. Its because the exercise price of 44 is below the average market price of $52. Hence, the treasury shares are repurchased at the average price for diluted EPS and is considered for computing exercise of options in the weighted average shares for Diluted EPS.

Basic EPS:

Net income = 3,050,000

Preferred Shareholder divided = 88,000

Basic EPS numerator = Net income - Preferred Shareholder divided

= (3,050,000- 88,000 ) = 2,962,000

Shares at Jan. 1 = 790,000

New Shares = 72,000

Adjustment as it was bought on February 28, = 72000 * 10/12 = 60,000

Shares retired = 6,000

Adjustment as it was reired on july 1, = 6000*6/12 = 3,000

Basic EPS denominator = (Shares at January 1 * Stock dividend adjustment) + (New Shares * Stock dividend adjustment) - Shares retired =  

=((790,000*(1+0.04))+((72,000*(10/12))*(1+0.04))-((6,000*(6/12))

= ((790,000*(1+0.04))+(60,000 *(1+0.04)) - 3,000

= 811,000

Basic EPS = 2,962,000/ 811,000 = 3.36

Diluted EPS:

Net income = 3,050,000

Preferred Shareholder divided = 88,000

After Tax Interest Saving = 500,000 * 10% = 50,000 - (50,000* 40%) = 30,000

Diluted EPS numerator = Net income - Preferred Shareholder divided + After Tax Interest Saving

= (3,050,000 - 88,000) +((500,000 * 10% ) - ((500,000 * 10%)*40%)))

= (3,050,000 - 88,000) + 30,000

= 2,992,000

Shares at Jan. 1 = 790,000

New Shares = 72,000

Adjustment as it was bought on February 28, = 72000 * 10/12 = 60,000

Shares retired = 6,000

Adjustment as it was reired on july 1, = 6000*6/12 = 3,000

Treasury shares are assumed repurchased at the average price for diluted EPS:

The options issued in 2016 are considered exercised for = 26,000 Shares

Exercise Price = 44

So, 26,000 shares * 44(exercise price) = 1,144,000

Average market price = 52

So, 1,144,000 / 52= 22,000 shares

Exercise of options = 26,000 Shares - 22,000 shares = 4,000

Contingent shares = 19,000 (shares already adjusted for the stock dividend)

Conversion of Bonds = (500,000  / 1000 ) * 40= 20,000

Diluted EPS denominator = (Shares at Jan * Stock Dividend Adjustment) + (New Shares * Stock Dividend Adjustment) - Shares retired + Exercise of Options + Contingent share + Conversion of Bonds =

= ((790,000 *(1+0.04))+((72,000 *(10/12)) * (1+0.04)) - ((6,000 *(6/12))) + (26,000 - 22,000 )+ 19,000 +((500,000/1000 )*40))

= ((790,000 *(1+0.04))+( 60,000* (1+0.04)) - 3000 + 4,000 + 19,000 + 20,000 =

= 924,000

Diluted EPS = 2,992,000/ 924,000= 3.24